Section 17  Control Theory and Optimization

Section 17: Control Theory and Optimization

Invited Lectures

Wednesday, 19 August  H 2032

16:15 – 17:00 Jorge Nocedal, Northwestern University, Evanston, USA Nonlinear Optimization: The Interplay Between Mathematical Characterizations and Algorithms (I: p.193) Chair: Margaret H. Wright

17:15 – 18:00 Jochem Zowe, Universität Erlangen, Erlangen, Germany What is the Best Material? — The Answer of a Mathematician (I: p.196) Chair: Josef Stoer

Thursday, 20 August  H 2032

14:00 – 14:45 Alexander Schrijver, CWI and University of Amsterdam, Amsterdam, Netherlands Routing and timetabling by topological search (I: p.194) Chair: Günter M. Ziegler

15:00 – 15:45 Michel X. Goemans, CORE, Louvain-La-Neuve, Belgium and MIT, Cambridge, USA Semidefinite programming and combinatorial optimization (I: p.193) Chair: András Frank

Friday, 21 August  H 2032

16:15 – 17:00 William J. Cook (also Section 16), Rice University, Houston, USA The Traveling Salesman Problem (I: p.186) Chair: Günter M. Ziegler

17:15 – 18:00 Jan C. Willems, University of Groningen, Groningen, The Netherlands Open Dynamical Systems and their Control (I: p.195) Chair: Fritz Colonius

Saturday, 22 August  H 2032

16:15 – 17:00 William R. Pulleyblank, IBM T. J. Watson Research Center, Yorktown Heights, USA Advances in Solving Airline Crew Pairing Optimization Problems (I: p.194) Chair: Rolf H. Möhring

Short Communications and Poster Sessions

Short Communications  Wednesday, 19 August
Applied Optimal Control  H 2033

Chair: Fredi Tröltzsch

14:00 – 14:15 Hans Josef Pesch, Technische Universität Clausthal, Clausthal, Germany Optimal Control of Large Scale DAE Systems in Chemical Engineering (p.347)

14:15 – 14:30 Fabian Flores-Bazan, Universitdad de Concepción, Concepcion, Chile Optimal solutions to some variational problems arising in Economics (p.341)

14:30 – 14:45 Bernd Kugelmann, Universität Greifswald, Greifswald, Deutschland Robust and Optimal Feedback-Control in Flight Dynamics (p.344)

Control Theory and Optimization  Section 17

Distributed Parameter Systems I  H 2033

Chair: Dan Tiba

15:00 – 15:15 Pierre, F. Aubert, INSA-Lyon, Vileurbanne, France Shape optimisation of geometrically nonlinear 3d beams and shells (p.336)

15:15 – 15:30 Raja Dziri, Faculté des Sciences de Tuni, Tunisi, Tunisia Dynamical shape optimazation in Navier-Stokes equations (p.339)

15:30 – 15:45 Michael Hinze, Technische Universität, Berlin, Germany On the construction of feedback controllers for parabolic equations with an application to the control of fluid flow (p.343)

Short Communications  Thursday, 20 August
Nonlinear Optimization I  H 2033

Chair: Jürgen Guddat

16:15 – 16:30 Leonid Faybusovich, University of Notre Dame, Notre Dame, IN, USA Interior-point algorithms and Euclidean Jordan algebras (p.341)

16:30 – 16:45 Bernd Kummer, Humboldt–Universität, Berlin, Germany Metric Regularity in Optimization: Trivial and Nontrivial Situations (p.344)

16:45 – 17:00 Ider Tsevendorj, Mongolian State University, Ulaanbaatar, Mongolia Sufficient Global Optimality Conditions for Nonconvex Nonsmooth Problems (p.353)

Convex Optimization  H 2033

Chair: Bernd Kummer

17:15 – 17:30 Anatoly Antipin, Russian Academy of Sciences, Moscow, Russia To Solving of Equilibrium Programming Problems (p.335)

17:30 – 17:45 Evgeny G. Belousov, Moscow State University, Moscow, Russia On Hausdorff continuity of convex polynomial mappings (p.336)

17:45 – 18:00 Patrick L. Combettes, City University, New York, NY, USA Some convergence principles in convex optimization (p.338)

Short Communications  Friday, 21 August
Optimality Conditions I  H 2033

Chair: Bernd Kugelmann

14:00 – 14:15 Vassili Dikoussar, Computing Centre RAS, Moscow, Russia Optimal Control with Control-State Constraints (p.339)

14:30 – 14:45 Olga O. Vasilieva, Universidad del Valle, Cali, Colombia Optimality Conditions of Singular Controls (p.354)

Variational Inequalities/Fractal Graphs  H 2013

Chair: Walter Gomez Bofill

14:00 – 14:15 Walter Gomez Bofill, Humboldt–Universität, Berlin, Germany On the regularity of one-parametric variational inequalities (p.341)

14:15 – 14:30 Olga A. Zinchenko, Karachai-Cherkes Technological Institute, Cherkessk, Russia The extreme graphs problems (p.356)

14:30 – 14:45 Anatolii Ivanovich Perov, Voronezh State University, Voronezh, Russia Discrete matrix Lur’e equation (singular case) (p.347)

Controllability and Nonlinear Systems  H 2033

Chair: Shanjian Tang

15:00 – 15:15 Serikbai A. Aisagaliev, Kazakh State University, Almaty, Kazakhstan Theory and constructive methods of solving the boundary problems of controllability processes

(p.335)

15:15 – 15:30 Zbigniew J. Bartosiewicz, Technical University of Bialystok, Bialystok, Poland Local observability, ideals, and finiteness conditions (p.336)

15:30 – 15:45 Shanjian Tang, Fudan University, Shanghai, China Brockett’s problem of classification of finite-dimensional estimation algebras for Nonlinear Filtering Systems (p.352)

Multiobjective Optimization  H 2013

Chair: Jan-J. Ruckmann

15:00 – 15:15 Jürgen Guddat, Humboldt–Universität, Berlin, Germany New Embeddings for Nonlinear Multiobjective Optimization Problems I (p.342)

15:15 – 15:30 Stefan M. Mititelu, Technical University of Civil, Bucharest, Romania Optimality and duality in nonsmooth vector programming (p.345)

Short Communications  Saturday, 22 August
Optimization in Function Spaces  H 2013

Chair: Valeri V. Obukhovskii

14:00 – 14:15 Andres, M. Encinas, UPC, Barcelona, Spain Existence and Approximation of Gauss Variational Problem in Potential Theory (p.340)

14:15 – 14:30 Stanisław Migórski, Jagiellonian University, Cracow, Poland Optimal control of hemivariational inequalities (p.345)

14:30 – 14:45 Valeri V. Obukhovskii, Voronezh State Pedagogical, Voronezh, Russia On some methods in a periodic problem for a functional differential inclusion in a Banach space

(p.346)

Stochastic Control/Controllability  H 2013

Chair: Galina A. Timofeeva

15:00 – 15:15 Yuri L. Sachkov, Program Systems Institute, Pereslavl-Zalessky, Russia Controllability of right-invariant systems on solvable Lie groups (p.349)

15:15 – 15:30 Leonid E. Shaikhet, Donetsk State University, Donetsk, Ukraine Stability of stochastic difference equations with varying delays (p.350)

15:30 – 15:45 Galina A. Timofeeva, Ural State Academy of Railway Transport, Ekaterinburg, Russia Effective control and observation in the multistage stochastic optimization problem (p.352)

Nonlinear Optimization II  H 2033

Chair: Leon A. Petrosjan

17:15 – 17:30 Alexei V. Blinov, Mari State University, Yoshkar-Ola, Russia The combined methods of centers with using reduced gradient for solving nonlinear optimization problems (p.337)

17:30 – 17:45 R. Enkhbat, Mongolian State University, Ulaanbaatar, Mongolia Global Optimality Conditions for some class of Nonconvex Optimization Problems (p.340)

17:45 – 18:00 Leon A. Petrosjan, St.-Petersburg State University, St.-Petersburg, Russia Time-consistancy problem in differential games (p.348)

Short Communications  Monday, 24 August
0/1 Problems  H 2033

Chair: Jannet A. Tambieva

16:15 – 16:30 Friedrich Eisenbrand, MPI Informatik, Saarbrücken, Germany On the Chvátal rank of polytopes in the 0/1 cube (p.339)

16:30 – 16:45 Anzor D. Kasaev, Karachai-Cherkes Technological Institute, Cherkessk, Russia Assigment Problem in Consideration of Discounting (p.344)

16:45 – 17:00 Jannet A. Tambieva, Karachai-Cherkes Technological Institute, Cherkessk, Russia Research of 4-Cliques Problem (p.351)

Algorithms and Complexity Issues  H 2033

Chair: Naum Zuselevich Shor

17:15 – 17:30 Vitaly A. Perepelitsa, Karachai-Cherkes Technological Institute, Cherkessk, Russia Fractal Graphs and their Properties (p.347)

17:30 – 17:45 Elena V. Popova, Karachai-Cherkes Technological Institute, Cherkessk, Russia Estimates of Complexity of the Vector Ordering Problem (p.348)

17:45 – 18:00 Naum Zuselevich Shor, Institute of Cybernetics, Kiev, Ukraine Methods of improving dual bounds with the use of functionally superfluous constraints in polynomial and boolean problems (p.350)

Short Communications  Tuesday, 25 August
Distributed Parameter Systems II  H 2033

Chair: Mikhail Zelikin

14:00 – 14:15 Andrzej Just, Technical University of Lodz, Lodz, Poland A Non-linear Hyperbolic Control Problem with Dirichlet Boundary Condition — Convergence of Galerkin Approximation (p.343)

14:15 – 14:30 Dan Tiba, Romanian Academy, Bucharest, Romania Recent Progress in the Analysis and Optimization of Plates (p.352)

14:30 – 14:45 Mikhail Zelikin, Department of Mathematics and Mechanics, Moscow, Russia Riccati-type PDE and Multidimensional Calculus of Variations (p.356)

Stochastic Programming I  H 2033

Chair: Werner Römisch

15:00 – 15:15 Vladimir Norkin, Glushkov Institute of Cybernetics, Kiev, Ukraine Stochastic optimization of discontinuous functions (p.346)

15:15 – 15:30 Nicole Groewe-Kuska, Humboldt–Universität, Berlin, Germany Confidence intervals for optimal values of simple recourse models (p.341)

15:30 – 15:45 Rüdiger Schultz, Mathematical Institute, Leipzig, Germany Mixed-Integer Linear Programming under Uncertainty (p.350)

Game Theory  H 2013

Chair: Hans Josef Pesch

16:15 – 16:30 Arkadii A. Chikrii, Cybernetics Institute, Kyiv, Ukraine

– to be announced – (p.338)

16:30 – 16:45 Greta Tsolakovna Chikrii, Cybernetics Institute, Nat. Ac. Sci. Ukraine, Kyiv, Ukraine Linear Differential Games with Variable Delay of Information (p.337)

16:45 – 17:00 Victor V. Zakharov, Saint-Petersburg University, Saint-Petersburg, Russia Subcore and reduced cooperative games (p.355)

Optimality Conditions II  H 2013

Chair: Henrikas Pragarauskas

17:15 – 17:30 Oleg I. Nikonov, Ural State Technical University, Ekaterinburg, Russia Methods of Guaranteed Control Theory in Finance (p.345)

17:30 – 17:45 Oleg Vladimirovich Vasiliev, Irkutsk State University, Irkutsk, Russian Federation High quality computational methods for optimal control problems (p.353)

17:45 – 18:00 Henrikas Pragarauskas, Institute of Mathematics and Informatics, Vilnius, Lithuania Integrodifferential Bellman equations (p.348)

Short Communications  Wednesday, 26 August
Stochastic Programming II  H 2033

Chair: Vladimir Norkin

14:00 – 14:15 René Henrion, WIAS Berlin, Berlin, Germany Quantitative stability of convex stochastic programs with probabilistic constraints (p.342)

14:15 – 14:30 Werner Römisch, Humboldt–Universität, Berlin, Germany Perturbation Analysis of Stochastic Programming Problems (p.349)

14:30 – 14:45 Roger Fandom Noubiap, Humboldt–Universität, Berlin, Germany A Minimax Algorithm for Calculating Statistical Tests under Generalized Moment Conditions

(p.340)

Miscellaneous I  H 2013

Chair: Mohamed Ali Hammami

14:00 – 14:15 Tomáš Roubíček, Charles University, Prague, Czech Republic Relaxation in optimization theory and variational calculus

14:15 – 14:30 Mohamed Ali Hammami, Faculté des Sciences de Sfax, Sfax-Tunisie, Tunisia Stabilization of a certain class of nonlinear systems via an observer

Semi-infinite and Semi-definite Optimization  H 2033

Chair: René Henrion

15:00 – 15:15 Jan-J. Ruckmann, Technische Universität München, München, Germany On Generalized Semi-infinite Optimization Problems (p.349)

15:15 – 15:30 Oliver Stein, Aachen University of Technology, Aachen, Germany On level sets of marginal functions (p.351)

15:30 – 15:45 Ivo Nowak, Humboldt–Universität, Berlin, Germany Bézier-Semidefinite Relaxations of nonconvex quadratic programs over a simplex (p.346)

Miscellaneous II  H 2013

Chair: Roland Wessäly

15:00 – 15:15 Ingo Bremer, WIAS, Berlin, Germany Experiences with the Optimization of Batch Distillation Processes

15:15 – 15:30 Rafail Gabasov, Byelorussian State University, Minsk, Belarus Synthesis of optimal dynamical systems

15:30 – 15:45 Roland Wessäly, Konrad-Zuse-Zentrum für Informationstechnik, Berlin, Germany New inequalities for a capacitated survivable network design problem (p.355)

Poster Session  Wednesday, 26 August 16:15 – 18:00  MA 144

Shamsha A. Aipanov, Kazakh State University, Almaty, Kazakhstan

The solutions of some problems of optimal control of spacecraft (p.335) Vladimir Borisov, Moscow State (Lomonosov) University, Cherkizovo, Russia

Bundles with chattering extremals (p.337)

Anna Debinska-Nagorska, Technical University of Lodz, Lodz, Poland A Non-linear Hyperbolic Control Problem with Neumann Boundary Condition — Convergence of Galerkin Approximation (p.338)

Roland Hildebrand, Universität Berlin, Berlin, Germany

Optimal synthesis of controlled systems near a focus (p.343) Christian Schulz, University of Dortmund, Dortmund, Germany

Optimal scheduling of a chemical batch plant (p.350) Zdzislaw Stempien, Technical University of Lodz, Lodz, Poland

On the Optimal Control of Nonlinear Hyperbolic System — Approximation (p.351) Pierre von Mouche, Landbouwuniversiteit Wageningen, Wageningen, The Netherlands

Sur deux problèmes géométriques dans la théorie des jeux (p.354) Olga Winzer, Zaporozhye State University, Ukraine

Complexity of vector spanning trees problems with topological criteria (p.355) Stanislav T. Zavalishchin, Institute of Mathematics and Mechanics, Ekaterinburg, Russia

Dynamic impulse systems. Optimization problems in fluid mechanics (p.356)